Required information

[The following information applies to the questions displayed below.]

Warnerwoods Company uses a perpetual inventory system. It entered into the following purchases and sales transactions for March.

Date Activities Units Acquired at Cost Units Sold at Retail
Mar. 1 Beginning inventory 240 units @ $53.80 per unit
Mar. 5 Purchase 295 units @ $58.80 per unit
Mar. 9 Sales 400 units @ $88.80 per unit
Mar. 18 Purchase 155 units @ $63.80 per unit
Mar. 25 Purchase 290 units @ $65.80 per unit
Mar. 29 Sales 270 units @ $98.80 per unit
Totals 980 units 670 units

4. Compute gross profit earned by the company for each of the four costing methods. For specific identification, the March 9 sale consisted of 135 units from beginning inventory and 265 units from the March 5 purchase; the March 29 sale consisted of 115 units from the March 18 purchase and 155 units from the March 25 purchase. (Round weighted average cost per unit to two decimals and final answers to nearest whole dollar.)
Answer #1

Solution:

Beginning inventory + purchases = 240 + 295 + 155 + 290 = 980units

sales = 400 + 270 = 670units

Ending inventory = 980 - 670 = 310units

Sales = ( 400 x 88.8) + (270 x 98.8) = 62,196

FIFO:

Cost of goods sold = (240 x 53.8) + (295 x 58.8) + (135* x 63.8) = 12,912 + 17,346 + 8,613 = 38,871

*135 = 670 - 240 - 295

Ending inventory = (20 x 63.8) + (290 x 65.8) = 1276 + 19082 = 20,358

155 - 135 = 20

Gross profit = sales - cost of goods sold = 62,196 - 38871 = 23,325

LIFO:

Cost of goods sold = (295 x 58.8) + ( 105* x 53.8) + (270 x 65.8) = 17346 + 5649 + 17766 = 40,761

105 = 400-295

Ending inventory = (135 x 53.8) + (155 x 63.8) + (20 x 65.8) = 7263 + 9889 + 1316 = 18,468

Gross profit = sales - cost of goods sold = 62196 - 40761 = 21,435

Weighted average method:

Weighted averge perpetual
Goods purchased Cost of goods sold Inventory Balance
Date # of units Cost per
unit
# of units sold Cost per
unit
Cost of
goods sold
# ofunits Cost per
unit
Inventory
Balance
Mar-01 240 @ 53.8 = 12912
Mar-05 295 @ 58.8 240 @ 53.8 = 12912
295 @ 58.8 = 17346
Average 535 @ 56.56 = 30258
Mar-09 400 @ 56.56 = 22624 135 @ 56.55 = 7634
Mar-18 155 @ 63.8 135 @ 56.55 = 7634
155 @ 63.8 = 9889
Average 290 @ 60.42 = 17523
Mar-25 290 @ 65.8 135 @ 56.55 = 7634
155 @ 63.8 = 9889
290 @ 65.8 = 19082
580 @ 63.11 = 36605
Mar-29 270 @ 63.11 = 17039.7 310 @ 63.11 = 19565.3
Totals 39663.7 19565.3

Gross profit = sales - cost of goods sold = 62,196 - 39664 = 22,532

Specific identification:

Cost of goods sold = (135 x 53.8 ) + (265 x 58.8 ) + (115 x 63.8 ) + (155 x 65.8) = 7263 + 15,582 + 7337 + 10,199 = 40,381

Ending inventory = (105 x 53.8) + (30 x 58.8) + (40 x 63.8) + (135 x 65.8) = 5649 + 1764 + 2552 + 8883 = 18,848

Gross profit = sales - cost of goods sold = 62,196 - 40381 = 21,815